Suppose that the production function of the firm is:
Q = 100L1/2.K1/2
K= 100, P = $1, w = $30 and r = $40. Determine the quantity of labor that the firm should hire
in order to maximize the profits. What is the maximum profit of this firm?
1
Expert's answer
2020-09-14T06:28:07-0400

The production function of the firm is given by;

"Q=100L ^{0.5}K^{0.5}"

Where Q is the total output produced by the firm, L is the amount of labor employed in the firm, K is the amount of capital employed in the firm and "k=100"


The cost function for the given firm is;

"C=wL+rK"

Where "w" is the wage rate, "\$30" and "r" is the interest rate "\$40" .

So now the cost function is given by;


"C= 30L+40K"

"C=30L+40(100)"

"C=30L+4000"

Where "K=100"

The market price is given as "p=\$1"


So the revenue function of the firm is given as;

"R=P\times Q"

"=1\times100L ^{0.5}(100)^{0.5}"

Put the value of 

"K(100)"

"R=100L^{0.5}(10)^{\frac{2}{2}}"


"\\implies R=100L \\times L^{0.5}\n\u200b"

So now the profit function of the firm is given by;


"\u03c0=R\u2212C"


Put the values of the revenue function (R) and the cost (c)

"\u03c0=100L^{0.5}\u221230L\u22124000"

Firm has to maximize its profit


To maximize profit, the First Order Condition (F.O.C) must be satisfied. For the F.O.C, take the differentiation with respect to 'L'


"\frac{δπ}{\delta L}=\frac{1000}{2}L^-{\frac{1}{2}}-30"


F.O.C: "\frac{δL}{δπ} =0"


"500L^-{\frac{1}{2}} −30=0"


"⟹\frac{500}{L^{0.5}}=30 ⟹30L^{0.5}=500"

Squaring both sides of the equation;


"900L=250,000\u27f9L=277.7777\u2248278"


The quantity of labor that the firm should hire in order to maximize the profits is

"L=278\ units"


The quantity of labor that the firm should hire in order to maximize the profits is

"L=278\ units"


"\pi =1,000L^{0.5}-30L-4,000"

put the value of "L=278"


"\pi = 139,000 - 8,340-4,000 \implies \pi = 126,660"

"\pi = 126,660"


The maximum profit of the firm is

"\pi = 126,660"